OpenStudy (anonymous):

Solve for x: 3 − (2x − 5) < −4(x + 2) (1 point) x < −8 x > −3 x > −8 x < −3

OpenStudy (anonymous):

dont you have to get all of the x-s on one side?

OpenStudy (anonymous):

3*5< -2x+-8 15<-2x-8 +8 +8 23<-2x 11.5>x...... uugh

OpenStudy (misty1212):

3 is not being multiplied by anything

OpenStudy (anonymous):

oops its + isnt it

OpenStudy (misty1212):

\[3 − (2x − 5)=3-2x+5=-2x+8\]

OpenStudy (misty1212):

start with \[-2x+8<-4x-8\]

OpenStudy (anonymous):

8<-2x+-8?

OpenStudy (misty1212):

add \(4x\) to both sides to get \[2x+8<-8\]

OpenStudy (anonymous):

and then - 8?

OpenStudy (misty1212):

subtract 8 get \[2x<-16\]

OpenStudy (misty1212):

then divide by 2 and you are done

OpenStudy (anonymous):

so x> -8

OpenStudy (anonymous):

oooh thank you!!